OpenStudy (anonymous):

write an equation in point slope for given: slope, m, = -1/5 and point (5,-2)

OpenStudy (anonymous):

Can you paste the point slope formula. Just google for the point slope formual, and i'll show you how to solve this.

OpenStudy (anonymous):

\[y+2=\frac{ -1 }{ 5 }(x-5)\]

OpenStudy (anonymous):

that is the point slope form

OpenStudy (anonymous):

@zordoloom point slope form: y-y1=m(x-x1)

OpenStudy (anonymous):

@dewsters Please don't just give out answers.